# A quantum implementation of high-order power method for estimating geometric entanglement of pure states

### v1

@@ 2.2.1 Recovering one-qubit states with tomography
From Eq. (7) it is evident that the components of |$\ket{{\bf v}_{i}^{(k)}}$ should be obtained under the condition that all the qubits except the $i$-th one are projected on the state |$\ket{0}$. Thus, to obtain the angles $(\vartheta_{i}^{(k)},\varphi_{i}^{(k)})$ for the one-qubit state encoding of |$\ket{{\bf v}_{i}^{(k)}}$ (see Eq. (5)) it is sufficient to solve the following system of equations:
$P_{i}(0)=\cos^{2}(\vartheta/2);$ (9)
$\langle X_{i}\rangle=P_{i}(-)-P_{i}(+)=\sin\vartheta\sin\varphi;$
$\langle Y_{i}\rangle=P_{i}(+i)-P_{i}(-i)=-\sin\vartheta\cos\varphi,$
where $X_{i}=\Pi_{0}^{\otimes(n-i)}\otimes X\otimes\Pi_{0}^{\otimes(i-1)}$ and $Y_{i}=\Pi_{0}^{\otimes(n-i)}\otimes Y\otimes\Pi_{0}^{\otimes(i-1)}$ with $\Pi_{0}=\ket{0}\bra{0}$ | being projector on the state |$\ket{0}$; the function
$P_{i}(s)=|\langle s\ket{{\bf v}_{i}^{(k)}}|^{2}=|v_{i,s}^{(k)}|^{2}=\frac{|u_{i,s}^{(k)}|^{2}}{\sum_{b\in B_{s}}|u_{i,b}^{(k)}|^{2}}$ (10)
is a probability of the $i$-th qubit being in the state |$\ket{s}=U_{s}\ket{0}\in B_{s}$ after the measurement, whilst all other qubits are in the state |$\ket{0}$, with $B_{s}=\{\ket{s},\ket{s}^{\perp}\}$ being a set of one-qubit orthonormal basis states and $u_{i,j}^{(k)}$ defined by Eq. (7).
[…]
Another method of measuring $u_{i,s}^{(k)}$ can be built upon the well-known Hadamard test procedure [41] as follows.
@@
    Introduce an ancilla qubit, initialized in the state |$\ket{+}=H\ket{0}$, whilst the working (data) qubits are initialized in the ground state |$\ket{\bf 0}$.
@@
    Perform a unitary operation $W={V_{i}^{(k)}}^{\dagger}U_{\psi}$ on the working qubits controlled by the ancilla qubit.
@@
    Perform a unitary operation $U_{s}^{\dagger}$ on the $i$-th qubit controlled by the ancilla qubit.
@@
    Measure $\langle X\rangle_{a}$ and $\langle Y\rangle_{a}$ on the ancilla qubit, and calculate $|u_{i,s}^{(k)}|^{2}=\langle X\rangle_{a}^{2}+\langle Y\rangle_{a}^{2}$.
The circuit representation of these steps is given in Fig. 2(b). This approach allows us to obtain each of the required probabilities by measuring six quantities, $P_{i}(s)$, for any number of qubits, however it requires one to be able to implement the controlled $U_{\psi}$ and $V_{i}$, which might be hard on the current generation of quantum devices for some cases of |$\ket{\psi}$ and large qubit systems.
In the following text we refer to this whole procedure as one-qubit tomography $\mathcal{T}_{i}$ for the $i$-th qubit on the given $n$-qubit state |$\ket{\mu}=W\ket{\bf 0}$:
$\mathcal{T}_{i}:$ $\mathbb{H}^{n}\to[0,\pi)\times[0,2\pi)$
|$\ket{\mu}\mapsto(\vartheta,\varphi).$
@@ 2.2.2 Measuring Entanglement
From this Remark we see that $\lambda^{(k)}$ is obtained by measuring $v_{i,0}^{(k)}$ using the procedure described in Section 2.2.1 with $W={V^{(k)}}^{\dagger}U_{\psi}$ (see Fig. 2(b)). We will denote this operation as
${\Lambda}:$ $\mathbb{H}^{n}\to[0,1]$
|$\ket{\mu}\mapsto\lambda.$
@@
Given a representation of a quantum circuit $U_{\psi}$ of a target state |$\ket{\psi}$ and a sufficiently small $\epsilon$, QHOPM (Algorithm 2) returns a pair $\lambda$ and $\{(\vartheta_{i},\varphi_{i})\}_{i=1\ldots n}$. Here $\{(\vartheta_{i},\varphi_{i})\}_{i=1\ldots n}$ are the parameters of the encoding of a separable state |$\ket{\phi}$. These values obey the following conditions:
@@
To show the other direction, having the components $v_{i,b}^{(k+1)}$, obtained in lines 4 and 6 of HOPM, we solve the system of equations (9) and recover the angles $(\vartheta_{i}^{(k+1)},\varphi_{i}^{(k+1)})$. This proves the equivalence of lines 4-6 of HOPM to line 3 of QHOPM. This result does not depend upon $i$ or $k$, thus, it is valid for any $i$ and $k$.
@@ 2.2.3 Algorithm Complexity
Recall that $n$ is the number of qubits in the input quantum circuit [^1] One may assume the circuit description of $U_{\psi}$ is of length polynomial in $n$. $U_{\psi}$ that defines the target state |$\ket{\psi}$. Assuming the use of the Hadamard test for the $n$-qubit target state |$\ket{\psi}$, QHOPM requires $n+1$ qubits.
Each application of the tomography step at line 3 of Algorithm 2 requires $12$ measurements (4 measurements each for $X$, $Y$, and $Z$); and each estimation of the scalar product at line 4 of Algorithm 2 requires $2$ measurements. Thus, each iteration of the loop in line 2 requires $212n+$ 2 measurements. Each measurement (assuming the Hadamard test) requires $\mathsf{O}\left(\epsilon^{-2}\right)$ single-shot readouts [^2] Note it is possible to reduce the amount of measurements required to $O(\epsilon^{-1})$ via amplitude amplification techniques. due to the Chernoff bound for absolute error $\epsilon$. The total shot complexity of each iteration is $\mathsf{O}\left(n\epsilon^{-2}\right)$.
The depth of the circuit used for measurements is defined by the depth of the circuit $U_{\psi}$ plus a constant. The constant is the sum of the depth of the separable state encoding ($V_{i}^{(k)}$ or $V^{(k)}$) and the overhead caused by the Hadamard test.
QHOPM is a direct quantum implementation of the classical HOPM algorithm and so has the same convergence properties in the complex domain [26, 29]. However, the rate of convergence has only been shown for real tensors [42, 43, 44]. To our knowledge there is no such analysis for the complex case.
@@ 2.3 Simulation Results
In this section, we simulate our algorithm to evaluate its robustness to noise using the IBM Qiskit [45] platform (see Methods section for simulation details). We also show how to mitigate the effects of noise on the algorithm. We refer to the target states by the notation State[$n$] for $n$ qubits, for example GHZ[$9$] is the GHZ state with 9 qubits and Random[$3$] is a particular random state with 3 qubits.
To demonstrate the convergence of the algorithm for different initial separable states we run each simulation with 10 random initial separable states. We then analyse $\overline{E}_{G}$, the average $E_{G}$ for the final 5 algorithm iterations over all initial separable states, and its standard deviation $\sigma_{E_{G}}$.
@@
We begin by verifying that our quantum implementation matches the classical algorithm without noise. We use the GHZ[$9$] state, which has a known $\widetilde{E}_{G}=0.50$, as an example. In Fig. 3(a) we see a good correspondence with classical HOPM, that is QHOPM converges to the same value ($\overline{E}_{G}\approx 0.50$, $\sigma_{E_{G}}\approx 2.8\times 10^{-4}$) as classical HOPM ($\overline{E}_{G}=0.50$). Here we are using $1\times 10^{7}$ shots per measurement yielding an expected error of $\epsilon\approx 3.2\times 10^{-4}$.
@@
In Fig. 3(b) we reduced the number of shots to $1\times 10^{5}$ and we see convergence to the same value $\overline{E}_{G}\approx 0.50$, but with slightly higher $\sigma_{E_{G}}\approx 8\times 10^{-3}$. Notice, that the distribution of $E_{G}$ values found after 1 iteration by QHOPM is different than that found by HOPM for the same initial separable states. By the second iteration HOPM and QHOPM converge to the same values. This divergence (from our experience) is normally not so noticeable, but of all the states we have studied (see Methods for the list of studied states), this effect is most pronounced in $GHZ[9]$. We assume that resolution errors, due to the reduced number of shots in the measurements, are amplified cumulatively for certain states (GHZ[$9$], in particular) during one-qubit tomography resulting in a different (but still improved) separable state from what HOPM or QHOPM with infinite resolution would find. From this point on we proceed with $1\times 10^{5}$ shots for all simulations, since it gives a sufficiently good convergence w.r.t. classical HOPM.
We then introduced depolarising noise (DN) (see Methods) with error rate $p=0.01$. The effects of DN are seen in Fig. 3(c) where the quantum algorithm still managed to converge with little variance ($\sigma_{E_{G}}\approx 0.009$), however, the value that it converged to ($\overline{E}_{G}\approx 0.678$) is greater than the true value ($\widetilde{E}_{G}=0.50$).
Figure 3: Simulation results for QHOPM approximating $E_{G}$ of $GHZ[9]$ in contrast to HOPM. (a) no noise with $1\times 10^{7}$ shots per measurement; (b) no noise with $1\times 10^{5}$ shots per measurement; (c) depolarizing noise, applied to all the gates with $p=0.01$. The mean geometric entanglement ($E_{G}$) for the fixed 10 random initial separable states per iteration of an algorithm is shown by lines. The shaded areas indicate the standard deviations. For the quantum simulations, individual values are marked with $\times$. All the simulations were run for $1\times 10^{5}$ shots per measurement.
@@
To explore the relationship between DN error rate $p$ and the divergence of QHOPM’s $E_{G}$ value from HOPM’s, we ran simulations with $p\in\{0.001,0.01,0.05\}$ representing acceptable, bad and terrible levels of noise, respectively. Again we used GHZ[$9$] as the target state. In Fig. 4(a) we see that when $p=0.001$ the convergence of QHOPM is only slightly off ($\overline{E}_{G}\approx 0.520,\sigma_{E_{G}}\approx 0.011$). The upper limit of error rate that is considered acceptable at this early stage of quantum computing is $p=0.01$, at this level the noise is causing our algorithm to converge to the value $\overline{E}_{G}\approx 0.678$ with $\sigma_{E_{G}}\approx 0.009$. At $p=0.05$ the algorithm still converges with similar variance as before ($\sigma_{E_{G}}\approx 0.005$) but to the incorrect value $\overline{E}_{G}\approx 0.946$.
We also examine the state Random[$6$] (see Methods) where we do not know the true $E_{G}$ value. Classical HOPM converges to the value $\overline{E}_{G}\approx 0.639$. With low noise ($p=0.001$) QHOPM converges to $\overline{E}_{G}\approx 0.659,\sigma_{E_{G}}\approx 0.012$, within $\epsilon\approx 0.02$ w.r.t. the classical algorithm’s result. Again, we see that as noise increases, the converged value increases: for $p=0.01$ it goes to $\overline{E}_{G}\approx 0.787$ with $\sigma_{E_{G}}\approx 0.011$, and for $p=0.05$ it goes to $\overline{E}_{G}\approx 0.978$ with $\sigma_{E_{G}}\approx 0.003$. We note the standard deviation of the converged values is within our measurement tolerance $\epsilon\approx 0.003$ (defined by the chosen number of shots; see Methods). We see the same effects for: GHZ[$3$], GHZ[$6$] (Supplemental Figure 6(a),(c)); Random[$3$], Random[$9$] (Supplemental Figure 9(a),(e)); the well studied W[$3$] state with $\widetilde{E}_{G}=5/9$ (Supplemental Figure 7(a)); and the important ring cluster states Ring[$3$], Ring[$6$], Ring[$9$], (Supplemental Figure 8(a),(c),(e)). We also observe that the shift in noise increases as the number of qubits increases.
We then characterised the shift in $\overline{E}_{G}$ induced by the noise as a function of circuit depth and noise rate. The resulting model (see Methods), based on the simple assumption that the depolarizing channel is applied with a constant error rate $p$ after each of $d$ layers of gates, yields a formula Eq. (25) that approximates the mitigated GE, $E_{G}^{\prime}$, expected without without noise.
We observe in Fig. 4(b) the effect of the mitigation on GHZ[$9$]. For $p=0.001$, the mitigated value $E_{G}^{\prime}\approx 0.500$, which corresponds to the true value. However for $p=0.01$, $E_{G}^{\prime}\approx 0.515$ is the result is slightly under-corrected. For $p=0.05$ mitigation is still helpful and adjusts the noisy value to $E_{G}^{\prime}\approx 0.558$. We see in Fig. 4(c) that the mitigation for the state Random[$6$] is extremely helpful: for $p=0.001$ $E_{G}$ being mitigated to $E_{G}^{\prime}\approx 0.640$, for $p=0.01$ to $E_{G}^{\prime}\approx 0.637$, and for $p=0.05$ to $E_{G}^{\prime}\approx 0.658$.
We suspect that the observed improvement in mitigation is due to the increased gate density of the Random[$6$] circuit (compared to GHZ[$9$]). Our model assumes the noise channel is applied to all qubits $d$ times (for circuit depth $d$) with the same error rate $p$. However, the DN implementation in Qiskit applies the channel only when there is a gate applied to the qubit. The structure of the circuit for GHZ[$9$] is one of the worst cases for our mitigation method since it includes a chain of 8 CNOTs traversing the 9 qubits one by one, leading to a deep but sparse circuit structure.
We also applied or mitigation technique to other GHZ states (Supplemental Fig. 6(b),(d),(f)), the W state (Supplemental Figure 7(b)), some Cluster Ring states (Supplemental Figure 8(b),(d),(f)), and some Random states (Supplemental Figure 9(b),(d),(f)).
@@
Figure 4: Simulation results for QHOPM on target states GHZ[$9$] and Random[$6$] with the DN model and its mitigation. (a) and (c) show the mean simulation results for QHOPM with input target states GHZ[$9$] and Random[$6$], respectively, for different DN error rates; (b) and (d) show the results of our mitigation procedure (see Eq. (25)) applied to QHOMP simulations on target statesGHZ[$9$] and Random[$6$], respectively. Each simulation was run 10 times (with the same 10 random initial separable states) with $1\times 10^{5}$ shots per each measurement. Solid lines represent the mean geometric entanglement, $E_{G}{}$, and the shaded colours represent the standard deviation. The colour of the lines represents the noise model used for the simulation (see legend). Red dotted lines show results of classical HOPM.
@@
Finally, we apply our migitation technique to results obtained with more realistic noise models. We consider the Qiskit “FakeLima” (representing an older 5 qubit device) and “FakeSherbrooke” (representing a newer device) models that are calibrated to match the error profile of physical devices. To apply our mitigation technique we need to estimate the noise rate $p$ for DN that would approximate the quantum device’s noise. To do this we apply QHOPM to the GHZ state, where the true value of $\widetilde{E}_{G}$ is known. We then use the ratio between the measured $\overline{E}_{G}$ and the true value $\widetilde{E}_{G}$ to estimate $p$ (see Methods, Eq. (26)).

### v2

@@ 2.2.1 Recovering one-qubit states with tomography
To obtain the angles $(\vartheta_{i}^{(k)},\varphi_{i}^{(k)})$ for the one-qubit state |$\ket{{\bf v}_{i}^{(k)}}$ of the $i$-th qubit at an arbitrary iteration $k$ of QHOPM, we use the following one-qubit tomography procedure. Let $\mathcal{T}_{i}\colon\mathrm{U}(2^{n})\to[0,\pi)\times[0,2\pi)$, which for some $W$ returns the angles $(\vartheta_{i},\varphi_{i})$, such that for |$\ket{w}=W\ket{{\bf 0}}$, |$\ket{q_{i}}=R_{z}(\varphi_{i})R_{x}(\vartheta_{i})\ket{0}$ and any one-qubit basis state |$\ket{s}$:
$\braket{s}{q_{i}}=\frac{\braket{s_{[i]}}{w}}{\sqrt{|\braket{s_{[i]}}{w}|^{2}+|\braket{s^{\perp}_{[i]}}{w}|^{2}}},$ (9)
and |$\ket{s^{\perp}_{[i]}}$ is an orthogonal state to |$\ket{s_{[i]}}$ defined in Eq.  (7).
In our implementation, $\mathcal{T}_{i}$ solves (using a classical device) the following system of equations
$\braket{Z}_{i}$ $=1-2P_{i}(0,W)=-\cos\vartheta_{i};$ (10)
$\braket{X}_{i}$ $=1-2P_{i}(+,W)=\sin\vartheta_{i}\sin\varphi_{i};$
$\braket{Y}_{i}$ $=1-2P_{i}(-i,W)=-\sin\vartheta_{i}\cos\varphi_{i},$
where $\braket{A}_{i}=\bra{q_{i}}A\ket{q_{i}}$; for the quantum system prepared in the state |$\ket{w}$,
$P_{i}(b,W)=|\braket{b}{q_{i}}|^{2}=\frac{|\braket{b_{[i]}}{w}|^{2}}{|\braket{b_{[i]}}{w}|^{2}+|\braket{b^{\perp}_{[i]}}{w}|^{2}}$ (11)
is a probability of the $i$-th qubit being in the state |$\ket{b}$, whilst other qubits are in the state |$\ket{0}$. The probabilities $P_{i}(b,W)$ are obtained by querying the quantum system.
One way to obtain $P_{i}(b,W)$ is by direct measurement of $|\langle b_{[i]}\ket{w}|^{2}$, however the number of shots for a given accuracy in this approach grows exponentially with the number of qubits. In our implementation, we use a Hadamard test-based [41] procedure where the number of shots depnds only on the accuracy and not the number of qubits.
First, we classically reconstruct the coefficients $C_{s}=\braket{s}{q_{i}}$ (defined by Eq.  (9)) for the state
|$\ket{q_{i}}=C_{b}\ket{b}+C_{b^{\perp}}\ket{b^{\perp}}$
in some basis $B_{b}=\{\ket{b},\ket{b^{\perp}}\}$. For this we measure the real and imaginary parts of $\braket{s_{[i]}}{w}$ for each basis state |$\ket{s}\in B_{b}$, using the Hadamard test procedure as follows.
@@
    Introduce an ancilla qubit, initialised in the state $\ket{+}$. The other qubits (referred to as data qubits in what follows) are initialised in the ground state |$\ket{\bf 0}$.
@@
    Perform a unitary operation $W$ on the data qubits, controlled by the ancilla qubit.
@@
    Perform a unitary operation $U_{s}^{\dagger}$, that transforms $\bra{0}$ | to $\bra{s}$ | state, on the $i$-th data qubit, controlled by the ancilla qubit.
@@
    Measure $x_{a}=\braket{X_{a}}$ and $y_{a}=\braket{Y_{a}}$ on the ancilla qubit ($A_{a}=A\otimes I^{n}$).
  5. 5.
    Calculate $\braket{s_{[i]}}{w}=x_{a}+\imath y_{a}$ ($\imath$ is an imaginary unit).
In the case of QHOPM, $W={V_{i}}^{(k)\dagger}U_{\psi}$ and |$\ket{q_{i}}=\ket{{\bf v}_{i}^{(k)}}$ as defined in Eq.  (6).
@@ 2.2.2 Measuring Entanglement
From this Remark we see that $\lambda^{(k)}$ can be obtained by performing the Hadamard test-based measurement procedure as described in the previous Subsection with $W={V^{(k)}}^{\dagger}U_{\psi}$ (see Fig. 2(b)). We will denote this operation as
${\Lambda}\colon$ $\mathbb{H}^{n}\to[0,1]$
|$\ket{w}\mapsto\lambda.$
@@
Given a unitary operator $U_{\psi}\in\mathrm{U}(2^{n})$, that prepares a target state |$\ket{\psi}=U_{\psi}\ket{\mathbf{0}}$ and a sufficiently small $\epsilon$, QHOPM (Algorithm 2) returns a pair $\lambda$ and $\{(\vartheta_{i},\varphi_{i})\}_{i=1\ldots n}$. Here $\{(\vartheta_{i},\varphi_{i})\}_{i=1\ldots n}$ are the parameters of the encoding of a separable state |$\ket{\phi}$. These values obey the following conditions:
@@
To show the other direction, having the components $v_{i,b}^{(k+1)}$, obtained in lines 4 and 6 of HOPM, we solve the system of equations Eq.  (10) and recover the angles $(\vartheta_{i}^{(k+1)},\varphi_{i}^{(k+1)})$. This proves the equivalence of lines 4-6 of HOPM to line 3 of QHOPM. This result does not depend upon $i$ or $k$, thus, it is valid for any $i$ and $k$.
@@ 2.2.3 Algorithm Complexity
Recall that $n$ is the number of qubits in the input unitary $U_{\psi}$ [^1] One may assume the circuit description of $U_{\psi}$ is of length polynomial in $n$. that defines the target state |$\ket{\psi}$. Assuming the use of the Hadamard test for the $n$-qubit target state |$\ket{\psi}$, QHOPM requires $n+1$ qubits.
Each application of the tomography step on line 3 of Algorithm 2 requires $4$ measurements to recover |$\ket{\mathbf{v}_{i}^{(k)}}$, and each estimation of the scalar product at line 4 of Algorithm 2 requires $2$ measurements. Thus, each iteration of the loop in line 2 uses $24n+$ 2 measurements. Each measurement (assuming the Hadamard test) requires $\mathsf{O}\left(\epsilon^{-2}\right)$ single-shot readouts [^2] Note it is possible to reduce the amount of measurements required to $O(\epsilon^{-1})$ via amplitude amplification techniques. due to the Chernoff bound for absolute error $\epsilon$. The total shot complexity of each iteration is $\mathsf{O}\left(n\epsilon^{-2}\right)$.
The depth of the circuit used for measurements is defined by the depth of the circuit for $U_{\psi}$ plus a constant. The constant is the sum of the depth of the separable state encoding ($V_{i}^{(k)}$ or $V^{(k)}$) and the overhead caused by the Hadamard test.
QHOPM is a direct quantum implementation of the classical HOPM algorithm and so has the same convergence properties in the complex domain [26, 29]. While it is known that HOPM converges with complex tensors, however, the rate of convergence has only been shown (to our knowledge)for real tensors [42, 43, 44].
@@ 2.3 Simulation Results
In this section, we simulate our algorithm to evaluate its robustness to noise using the IBM Qiskit [45] platform (see Methods section for simulation details). We also show how to mitigate the effects of noise on the algorithm. We refer to target states by the notation State[$n$] for $n$ qubits, for example GHZ[$9$] is the GHZ state with 9 qubits and Random[$3$] is a particular random state with 3 qubits.
When using HOPM we choose the minimum $E_{G}$ from a sample of initial separable states, however (due to the finite number of shots) QHOPM tends to oscillate above and below the convergence value. We find the median gives a more accurate summary of the convergence value of QHOPM than the mean, perhaps reflecting some skew in the distribution.
For a given target state let $\mathcal{F}_{k}$ represent a set of estimations of $E_{G}$ for a sample of initial separable states at iteration $0\leq k\leq K$. Let $\overline{E}_{G}$ represent the value converged to by QHOPM, computed as the median of the values in =$\mathcal{F}=\{\text{m}(\mathcal{F}_{k})\mid k\in\left[K-5,\ldots K\right]\}$, where $\text{m}(\mathcal{A})$ is the median of the values in the set $\mathcal{A}$. Let $\mathrm{IQR}_{\overline{E}_{G}}$ be the interquartile range for $\mathcal{F}$. We note that the choice of 5 final iterations here is arbitrary (statistically better results are obtained if one uses more iterations). Here we choose the median instead of the mean to avoid possible outliers induced by certain initial separable states.
@@
We begin by verifying that our quantum implementation matches the classical algorithm and demonstrate the effect on convergence of the depolarising noise (DN) model (see Methods) with noise rate $p=0.01$. We use the GHZ[$9$] state, which has a known $\hat{E}_{G}=0.50$, as an example. In Fig. 3 we see a good correspondence with classical HOPM for $1\times 10^{7}$ shots per measurement (yielding an expected error of $\epsilon\approx 3.2\times 10^{-4}$ according to the Chernoff bound), that is QHOPM converges to the same value ($\overline{E}_{G}\approx 0.50$, $\mathrm{IQR}_{\overline{E}_{G}}\lesssim 10^{-4}$) as classical HOPM ($\overline{E}_{G}=\hat{E}_{G}=0.50$). Then we reduced the number of shots to $1\times 10^{5}$ ($\epsilon\approx 3.2\times 10^{-3}$) and we see convergence to $\overline{E}_{G}\approx 0.503$, within the slightly higher $\mathrm{IQR}_{\overline{E}_{G}}\approx 0.005$ of the true value. These results are consistent with the Chernoff bound discussed in the previous Section. Finally, we introduced depolarising noise (DN) (see Methods) with error rate $p=0.01$. Remarkably, the quantum algorithm still managed to converge with little variability ($\mathrm{IQR}_{\overline{E}_{G}}\approx 0.003$), however, the value it converged to ($\overline{E}_{G}\approx 0.679$) is greater than the true value.
Figure 3: The results of HOPM and QHOPM for (a) approximating the geometric entanglement $E_{G}$ of GHZ[$9$] and (b) the absolute error ($|\hat{E}_{G}-E_{G}|$). The lines represent the median value for 10 random initial separable states for different numbers of shots per measurement and noise models (see legend). The shaded areas show the interquartile ranges.
@@
Notice that the distribution of $E_{G}$ values found after 1 iteration by QHOPM is different from that found by HOPM for the same initial separable states. By the second iteration, HOPM and QHOPM converge to the same values. This divergence (from our experience) is normally not so noticeable, but of all the states we have studied (see Methods for the list of studied states), this effect is most pronounced in GHZ[$9$]. We assume that resolution errors, due to the reduced number of shots in the measurements, are amplified cumulatively for certain states (GHZ[$9$], in particular) during one-qubit tomography resulting in a different (but still improved) separable state from what HOPM or QHOPM with infinite resolution would find. From this point on we proceed with $1\times 10^{5}$ shots and the median over 10 initial separable states for all simulations (unless said otherwise) since it gives a sufficiently good convergence w.r.t. classical HOPM.
To evaluate the performance of QHOPM for random target states we ran QHOPM with no noise for 100 Random[$n$] target states each with 20 initial separable states with $n\in\{3,\ldots,6\}$ qubits. In Fig. 4 we summarize the results of the simulation. Here for each target state at iteration $k$ we calculate the median result $\text{m}(\mathcal{F}_{k})$ over 20 initial separable states and plot the median absolute difference $|\hat{E}_{G}-\text{m}(\mathcal{F}_{k})|$ (and the interquartile range) over all the target states for each $k$ and number of qubits $n$. We show that the median value of the absolute error $|\hat{E}_{G}-\text{m}(\mathcal{F}_{k})|$ converges to the values less than $10^{-2}$ for any number of qubits, which is again consistent with the Chernoff bound. In the case of negligible noise, the accuracy of QHOPM depends only on the number of shots per measurement.
Figure 4: The absolute error ($|\hat{E}_{G}-\text{m}(\mathcal{F}_{k})|$) of QHOPM approximating $E_{G}$ of Random[$n$] states with no noise. The lines represent the median absolute error over 100 Random[$n$] target states ($n\in\{3,\ldots,6\}$ qubits) between the best (minimal) HOPM results and the median results $\text{m}(\mathcal{F}_{k})$ computed by QHOPM for each target state over 20 random initial separable states. The shaded areas show the interquartile ranges for the absolute error.
@@
To explore the relationship between DN error rate $p$ and the divergence of QHOPM’s $E_{G}$ value from HOPM’s, we ran simulations with $p\in\{0.001,0.01,0.05\}$ representing “acceptable”, “bad” and “terrible” levels of noise, respectively. We used GHZ[$9$] and Random[$6$] as the target states. In Fig. 5(a) we see that for GHZ[$9$] when $p=0.001$ QHOPM converges to $\overline{E}_{G}\approx 0.519,\mathrm{IQR}_{\overline{E}_{G}}\approx 0.003$. The upper limit of error rate $p$ that is considered acceptable at this early stage of quantum computing is $0.01$, at this level, the noise is causing our algorithm to converge to the value $\overline{E}_{G}\approx 0.679$ with $\mathrm{IQR}_{\overline{E}_{G}}\approx 0.003$. At $p=0.05$ the algorithm still converges with a similar variability but to the incorrect value $\overline{E}_{G}\approx 0.946$ with $\mathrm{IQR}_{\overline{E}_{G}}\approx 0.002$. The results are also summarised in Table 1.
@@
Figure 5: Simulation results for QHOPM on target states GHZ[$9$] and Random[$6$] with the DN model and its mitigation. (a) and (c) show the simulation results for QHOPM with input target states GHZ[$9$] and Random[$6$], respectively, for different DN error rates; (b) and (d) show the results of our mitigation procedure (see Eq.  (26)) applied to results in (a) and (c), respectively. Solid lines represent the median geometric entanglement, $E_{G}{}$, over the results for different initial separable states and the shaded colours represent the interquartile range. The colour of the lines represents the DN noise rate used for the simulation (see legend). Red dotted lines show the results of classical HOPM.
Table 1: Results of HOPM and QHOPM with different DN noise rates $p$ for GHZ[$9$] and Random[$6$] states (see. Fig. 5). Minimum HOPM estimate for Random[$6$] is $E_{G}=0.636$.
Target DN $p$ Measured Mitigated
$\overline{E}_{G}$ $\mathrm{IQR}_{\overline{E}_{G}}$ $\overline{E}_{G}^{\prime}$ $\mathrm{IQR}_{\overline{E}_{G}^{\prime}}$
GHZ[$9$] 0.001 0.519 0.003 0.499 0.003
0.010 0.679 0.003 0.516 0.006
0.050 0.946 0.002 0.554 0.019
Random[$6$] 0.001 0.657 0.003 0.638 0.003
0.010 0.789 0.004 0.639 0.008
0.050 0.978 0.000 0.661 0.029
In Fig. 5(c) we give the results for the median $E_{G}$ for Random[$6$]. With low noise ($p=0.001$) QHOPM converges to $\overline{E}_{G}\approx 0.657,\mathrm{IQR}_{\overline{E}_{G}}\approx 0.003$, which is within $\epsilon\approx 0.02$ w.r.t. the classical algorithm’s result. Again, we see that as noise increases, the converged value increases: for $p=0.01$ it goes to $\overline{E}_{G}\approx 0.789$ with $\mathrm{IQR}_{\overline{E}_{G}}\approx 0.004$, and for $p=0.05$ it goes to $\overline{E}_{G}\approx 0.978$ with $\mathrm{IQR}_{\overline{E}_{G}}\approx 0.000$. We note that the interquartile range of the converged values is within our measurement tolerance $\epsilon\approx 0.003$ (defined by the chosen number of shots; see Methods).
We see the same effects for: GHZ[$3$], GHZ[$6$] (Supplemental Figure 7(a),(c)); Random[$3$], Random[$9$] (Supplemental Figure 10(a),(e)); the well studied W[$3$] state with $\hat{E}_{G}=5/9$ (Supplemental Figure 8(a)); and the important ring cluster states Ring[$3$], Ring[$6$], Ring[$9$], (Supplemental Figure 9(a),(c),(e)). We also observe that the shift, induced by noise, increases as the number of qubits increases.
We then characterised the shift in $\overline{E}_{G}$ induced by the noise as a function of circuit depth and noise rate. The resulting model (see Methods), based on the simple assumption that the depolarising channel is applied with a constant error rate $p$ after each of $d$ layers of gates, yields a formula Eq.  (26) that approximates the mitigated GE, $E_{G}^{\prime}$, expected without noise. Similarly to $\overline{E}_{G}$ we define $\overline{E}_{G}^{\prime}$ as the mitigated result of QHOPM, computed as the median of the values in =$\mathcal{F}^{\prime}=\{\text{m}(\mathcal{F}^{\prime}_{k})\mid k\in\left[K-5,\ldots K\right]\}$, where $\mathcal{F}^{\prime}_{k}$ represents a set of mitigated values $E^{\prime}_{G}$ for a sample of initial separable states at iteration $0\leq k\leq K$. Similarly $\mathrm{IQR}_{\overline{E}_{G}^{\prime}}$ is the interquartile range of $\mathcal{F}^{\prime}$. We observe in Fig. 5(b) the effect of the mitigation on GHZ[$9$]. For $p=0.001$, the mitigated value $\overline{E}_{G}^{\prime}\approx 0.499$ with $\mathrm{IQR}_{\overline{E}_{G}^{\prime}}=0.003$, which corresponds to the true value. However, for $p=0.01$, $\overline{E}_{G}^{\prime}\approx 0.516$ with $\mathrm{IQR}_{\overline{E}_{G}^{\prime}}=0.006$ is the result. For $p=0.05$ mitigation is still helpful and adjusts the noisy value to $\overline{E}_{G}^{\prime}\approx 0.554$ ($\mathrm{IQR}_{\overline{E}_{G}^{\prime}}=0.019$). We see in Fig. 5(d) that the mitigation for the state Random[$6$] gives better results: for $p=0.001$ $E_{G}$ being mitigated to $\overline{E}_{G}^{\prime}\approx 0.638$ ($\mathrm{IQR}_{\overline{E}_{G}^{\prime}}=0.003$), for $p=0.01$ to $\overline{E}_{G}^{\prime}\approx 0.639$ ($\mathrm{IQR}_{\overline{E}_{G}^{\prime}}=0.008$), and for $p=0.05$ to $\overline{E}_{G}^{\prime}\approx 0.661$ ($\mathrm{IQR}_{\overline{E}_{G}^{\prime}}=0.029$). We suspect that the observed improvement in mitigation is due to the increased gate density of the Random[$6$] circuit compared to GHZ[$9$]. Our model assumes the noise channel is applied to all qubits $d$ times (for circuit depth $d$) with the same error rate $p$. However, the DN implementation in Qiskit applies the channel only when there is a gate applied to the qubit. The structure of the circuit for GHZ[$9$] is one of the worst cases for our mitigation method since it includes a chain of 8 CNOTs traversing the 9 qubits one by one, leading to a deep but sparse circuit structure.
We also applied our mitigation technique to other GHZ states (Supplemental Fig. 7(b),(d),(f)), the W state (Supplemental Figure 8(b)), some Cluster Ring states (Supplemental Figure 9(b),(d),(f)), and some Random states (Supplemental Figure 10(b),(d),(f)).
@@
Finally, we apply our mitigation technique to results obtained with more realistic noise models. We consider the Qiskit “FakeLima” (representing an older 5 qubit device) and “FakeSherbrooke” (representing a newer device) Qiskit backends that are calibrated to match the error profile of physical devices. To apply our mitigation technique we need to estimate the noise rate $p$ for DN that would approximate the quantum device’s noise. To do this we apply QHOPM to the GHZ state, where the true value $\hat{E}_{G}$ is known. We then use the ratio between the measured $\overline{E}_{G}$ and the true value $\hat{E}_{G}$ to estimate $p$ (see Methods, Eq.  (27)).